# Effective synchronization amid noise-induced chaos

### v2

@@ A.2 Identifying a fiducial phase
In Sec. IV of the main text, we describe the estimation of a smooth probability function $Q(\varphi,\sigma)$ from a set of samples $\{\varphi_{n}\}$. In Fig. S7 we show the effect of a different bandwidth $\sigma$ using sampled $\{\varphi_{n}\}$’s from the main text.
Appendix S7 fiducial phase differences between agents
@@
In Sec. IV of the main text, we compared the fiducial phases obtained by two independent agents. Their oscillators were subjected to the same phase map $\psi(\varphi)$ and the same kick timings $\{\beta^{(k)}\}$. Since each agent used a finite set of $N=500$ samples to obtain their $\varphi_{\mathrm{f}}^{(k)}$ after the $k$’th kick, their fiducial phases may not fully coincide when they sample an identical distribution. Thus, the agents obtain different values for $\varphi_{\mathrm{f}}^{(k)}$. Figure S8 shows the obtained differences $|\Delta\varphi_{\mathrm{f}}|$ as a scatter plot against the common distribution’s entropy, using two different bandwidths, (a) $\sigma=e^{S}$ and (b) $\sigma=e^{S}/N$. We interpret these results in the next section.
Appendix S8 Rare misidentifications of fiducial phase
@@
We point out that out of $10^{4}$ data-points for $|\Delta\varphi_{\mathrm{f}}|$ used in Sec. IV of the main text, $88$ points lie above the $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ line ($0.9\%$), typically even a few orders of magnitude above it for high-entropy distributions; see Fig. S8(a). On examination, these points either corresponded to distributions with entropy $S\sim 0$ or proved to be misidentifications, in which the $Q(\varphi;e^{S})$’s had two peaks of nearly equal height and the agents made opposite determinations of the global maximum, as a result of the finite samples used to estimate $Q(\varphi;e^{S})$. Among these occurrences is the distribution of Fig. S7, where each agent chose a different peak out of the two at $0.41$ and $0.93$.
@@
These misidentifications were an important factor in our choice of smoothing width $\sigma=e^{S}$. Indeed, when we used $\sigma=e^{S}/N$ (rather than $\sigma=e^{S}$ of the main text) as a bandwidth, there were $791$ misidentified points ($8\%$), which have occurred with distributions of entropies as a low as S=14S=-14; see Fig. S8(b). In fact, the large concentration of points surrounding $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ suggests that the bandwidth is too small, and not enough samples are involved in estimating the peak center, raising the error from $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$ to just the overall “width” of a given distribution $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}$. At the same time, in the cases where the identification of the peak was precise (that is, the highest peak clearly dominates over the rest and encapsulated $\sim N$ samples), this smaller bandwidth allowed for an appropriately smaller discrepancy, $|\Delta\varphi_{\mathrm{f}}|\sim(e^{S}/N)/\sqrt{N}$. Thus, as we expected, smaller bandwidths facilitate smaller deviations, which comes at a cost of more frequent, finite-sample-size-related misidentifications of the highest peak.
Figure S1: Relation between the fiducial phase position $\varphi_{\mathrm{f}}$ and the instantaneous fiducial phase $\varphi_{\mathrm{f}}+0.2T$ at a later time $t=0.2T$, as measured by a given agent. Orange dotted curve represents the distribution $q(\varphi)$ for some particular noise history. The highest peak of this distribution is then $\varphi_{\mathrm{f}}$, and is marked with a dashed line. At time $t=0$ this is the actual phase distribution. At a time $t=0.2T$ later, every phase point, including the agent’s own oscilllator’s phase, has advanced through $0.2$ cycles. The probability distribution $q(\varphi-0.2T)$ and fiducial phase $\varphi_{\mathrm{f}}+0.2T$ at that moment is shown by the blue line.
Figure S2: Trajectories from the dynamical system of Eq. (S1). The limit cycle (period $T=1.98$) is shown in full black oval. The black ticks along it are equispaced $\varphi\in[0,1)$ values, separated by $\Delta t/T=0.05$. ($\varphi=0$ is located at $0+0.4i$, overlapping with the empty square.) We show the resulting trajectories from kicking oscillators positioned at $\varphi=0$ (empty square), $\varphi\simeq 0.35$ (empty triangle), and $\varphi\simeq 0.6$ (empty diamond) by (a) $z\to z+0.15$ and (b) $z\to z+0.45$. The first three small points on each trajectory are the positions of the oscillators after $t=0,(1/2)T,T$. After a time $9T$, the oscillators reach the stable orbit with new phase values, $\psi\simeq-0.13$ (full square), $\psi\simeq 0.73$ (full triangle), and $\varphi\simeq 0.36$ (full diamond) in (a), and $\psi\simeq-0.44$ (full square), $\psi\simeq 1.37$ (full triangle), and $\varphi\simeq-0.66$ (full diamond) in (b).
Figure S3: Phase lags from the dynamical system of Eq. (S1). Empty blue symbols are phase shifts as seen in Fig. S2. The square, diamond, and triangle symbols correspond to the three trajectories in Fig. S2. Since only the fractional parts of phases are significant, each symbol was shifted upwards or downwards by an integer so as to obtain a continuous dependence in $\varphi$. In addition, this curve would be shifted vertically and horizontally by choosing a different point on the orbit as the phase origin. Red points show the standardized phase map as used in the main text. Namely, the phase origin and integer-valued shifting were chosen so that $\psi(0)=0$. Further, each $\psi(\varphi)$ value was shifted to lie between 0 and 1.
Figure S4: Two phase distributions $q^{(k)}(\varphi)$ before (blue full circles) and after (orange empty circles, with inflected axes) a kick, obtained for a typical realization of the cubic phase map (Eqs. (1) and (2) of the main text) with $A=9.32$ ($\Lambda=0.141$). The cubic phase map is drawn with a red line. One sees how the ‘sporadic’ region of the initial distribution in the vicinity the phase map’s extremum ($\varphi\simeq 0.7$) is condensed into a prominent peak at $\psi\simeq 0.3$. Conversely, the surroundings of the peak at $\varphi\simeq 0.95$ are widened at $\psi=0.7$, while its height decreased from $\mathcal{O}(10)$ to $\mathcal{O}(1)$. The phase distributions are sampled from $N=500$ oscillators.
$\Lambda=-0.141$ Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption $\Lambda=0.141$ Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
Figure S5: A sequence of two, initially-identical phase distributions $q^{(k)}_{\mathrm{A}}(\varphi)$ (red) and $q^{(k)}_{\mathrm{B}}(\varphi)$ (blue) subjected to different noises in a typical realization of Eq. (1) and Eq. (2) of the main text with either $\Lambda=-0.141$ ($A=7.32$) or $\Lambda=0.141$ ($A=9.32$), as indicated. The initial distributions are uniform, $\varphi\sim\mathrm{U}[0.5,0.55)$. The distributions are drawn on a circle so the periodicity $\mathrm{mod}\,1$ of the phase circle is apparent. The radial axis shows the distributions’ values $q(\varphi)$ on a log scale, where $q(\varphi)=1$ for the inner full circle and as indicated for the outer dotted circles. The kick numbers $k$ are shown inside the inner circle. Since the noise is different for each ensemble, $\Lambda<0$ produces different synchronized states, whereas $\Lambda>0$ produces erratic distributions that do not converge unto each other. The distributions are sampled using $N=500$ oscillators via the nearest-neighbor distances; see Eq. (A1) of the main text. The initial samples drawn for ensemble $\mathrm{A}$} are copied to ensemble $\mathrm{B}$}, so the starting point is truly identical.
Figure S6: (a) Kullback-Leibler divergences (KLDs) $D(\mathrm{A}\|\mathrm{B})$ and $D(\mathrm{B}\|\mathrm{A})$ among two distributions and (b) entropies $S_{\mathrm{A}}$ and $S_{\mathrm{B}}$ of each distribution, as obtained for the waiting time realization $\{\beta^{k}\}$ of Fig. 2 of the main text with $\Lambda=-0.141$. The number of phase samples is $N=500$. The initial distribution per ensemble is obtained with width $u=0.05$. The numerical errors in the estimation of the KLDs and entropies are $\Delta D=0.05$ and $\Delta S=0.2$, respectively.
Figure S7: Illustration of the kernel-density estimation method. Here we estimate the distribution from $N=500$ phase samples obtained at $k=100$ by agent $\mathrm{B}$ during the realization of Fig. 2 of the main text for $\Lambda=0.141$. The density estimate of Eq. (A1) of the main text is shown in points. The kernel-density estimate, $Q(\varphi;\sigma)$, is carried out with either $\sigma=e^{S}$ (the bandwidth we use throughout Sec. IV of the main text; dashed red curve) or $\sigma=e^{S}/N$ (the alternative extreme; thin solid black curve). S=2.29S=-2.29 for this distribution using Eq. (A2) of the main text. A scale bar of $e^{S}$ is shown for reference with a thick black horizontal line. Clearly, the result of the latter bandwidth is not statistically significant, as it is comparable to the nearest-sample distances. The analysis in the main text is carried out with the former bandwidth.
Refer to captionRefer to caption
Figure S8: A scatter plot of deviations in the fiducial phases of each party $\Delta\varphi_{\mathrm{f}}^{(k)}=\varphi_{\mathrm{f},\mathrm{A}}^{(k)}-\varphi_{\mathrm{f},\mathrm{B}}^{(k)}$ against the entropy of their distributions $S^{(k)}=(S^{(k)}_{\mathrm{A}}+S^{(k)}_{\mathrm{B}})/2$, using bandwidths (a) $\sigma=e^{S}$ (used for Fig. 5 of the main text) and (b) $\sigma=e^{S}/N$. Each $(S,|\Delta\varphi_{\mathrm{f}}|)$ point is computed given a pair of distributions $Q_{\mathrm{A}}^{(k)}$ and $Q_{\mathrm{B}}^{(k)}$ obtained after kicks (with $\Lambda=0.141$) $k=1,\ldots 10^{4}$. For completeness, the points’ shading encodes the kick at which it was obtained, as indicated. The orange dotted, solid, and dashed lines are, respectively, $|\Delta\varphi_{\mathrm{f}}|=e^{S}$, $|\Delta\varphi_{\mathrm{f}}|=e^{S}/\sqrt{N}$, and $|\Delta\varphi_{\mathrm{f}}|=e^{S}/N^{3/2}$.

### v3

@@ Appendix H Effect of the bandwidth
In Sec. IV, we describe the estimation of a smooth probability function $Q(\varphi,\sigma)$ from a set of samples $\{\varphi_{n}\}$. In Fig. 12 we show the effect of a different bandwidth $\sigma$ using sampled $\{\varphi_{n}\}$’s.
Appendix I fiducial phase differences between agents
@@
In Sec. IV, we compared the fiducial phases obtained by two independent agents. Their oscillators were subjected to the same phase map $\psi(\varphi)$ and the same kick timings $\{\beta^{(k)}\}$. Since each agent used a finite set of $N=500$ samples to obtain their $\varphi_{\mathrm{f}}^{(k)}$ after the $k$’th kick, their fiducial phases may not fully coincide when they sample an identical distribution. Thus, the agents obtain different values for $\varphi_{\mathrm{f}}^{(k)}$. Figure 13 shows the obtained differences $|\Delta\varphi_{\mathrm{f}}|$ as a scatter plot against the common distribution’s entropy, using two different bandwidths, (a) $\sigma=e^{S}$ and (b) $\sigma=e^{S}/N$. We interpret these results in the next section.
Figure 12: Illustration of the kernel-density estimation method. Here we estimate the distribution from $N=500$ phase samples obtained at $k=100$ by agent $\mathrm{B}$ during the realization of Fig. 2 for $\Lambda=0.141$. The density estimate of Eq. (9b) is shown in points. The kernel-density estimate, $Q(\varphi;\sigma)$, is carried out with either $\sigma=e^{S}$ (the bandwidth we use throughout Sec. IV; dashed red curve) or $\sigma=e^{S}/N$ (the alternative extreme; thin solid black curve). S=2.29S=-2.29 for this distribution using Eq. (10). A scale bar of $e^{S}$ is shown for reference with a thick black horizontal line. Clearly, the result of the latter bandwidth is not statistically significant, as it is comparable to the nearest-sample distances. The analysis in Sec. IV.1 is carried out with the former bandwidth.
Appendix J Rare misidentifications of fiducial phase
@@
We point out that out of $10^{4}$ data-points for $|\Delta\varphi_{\mathrm{f}}|$ used in Sec. IV, $88$ points lie above the $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ line ($0.9\%$), typically even a few orders of magnitude above it for high-entropy distributions; see Fig. 13(a). On examination, these points either corresponded to distributions with entropy $S\sim 0$ or proved to be misidentifications, in which the $Q(\varphi;e^{S})$’s had two peaks of nearly equal height and the agents made opposite determinations of the global maximum, as a result of the finite samples used to estimate $Q(\varphi;e^{S})$. Among these occurrences is the distribution of Fig. 12, where each agent chose a different peak out of the two at $0.41$ and $0.93$.
Refer to captionRefer to caption
Figure 13: A scatter plot of deviations in the fiducial phases of each party $\Delta\varphi_{\mathrm{f}}^{(k)}=\varphi_{\mathrm{f},\mathrm{A}}^{(k)}-\varphi_{\mathrm{f},\mathrm{B}}^{(k)}$ against the entropy of their distributions $S^{(k)}=(S^{(k)}_{\mathrm{A}}+S^{(k)}_{\mathrm{B}})/2$, using bandwidths (a) $\sigma=e^{S}$ (used in Fig. 5) and (b) $\sigma=e^{S}/N$. Each $(S,|\Delta\varphi_{\mathrm{f}}|)$ point is computed given a pair of distributions $Q_{\mathrm{A}}^{(k)}$ and $Q_{\mathrm{B}}^{(k)}$ obtained after kicks (with $\Lambda=0.141$) $k=1,\ldots 10^{4}$. For completeness, the points’ shading encodes the kick at which it was obtained, as indicated. The orange dotted, solid, and dashed lines are, respectively, $|\Delta\varphi_{\mathrm{f}}|=e^{S}$, $|\Delta\varphi_{\mathrm{f}}|=e^{S}/\sqrt{N}$, and $|\Delta\varphi_{\mathrm{f}}|=e^{S}/N^{3/2}$.
@@
These misidentifications were an important factor in our choice of smoothing width $\sigma=e^{S}$. Indeed, when we used $\sigma=e^{S}/N$ (rather than $\sigma=e^{S}$ of Sec. IV.1) as a bandwidth, there were $791$ misidentified points ($8\%$), which have occurred with distributions of entropies as a low as S=14S=-14; see Fig. 13(b). In fact, the large concentration of points surrounding $|\Delta\varphi_{\mathrm{f}}|=e^{S}$ suggests that the bandwidth is too small, and not enough samples are involved in estimating the peak center, raising the error from $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}/\sqrt{N}$ to just the overall “width” of a given distribution $|\Delta\varphi_{\mathrm{f}}|\sim e^{S}$. At the same time, in the cases where the identification of the peak was precise (that is, the highest peak clearly dominates over the rest and encapsulated $\sim N$ samples), this smaller bandwidth allowed for an appropriately smaller discrepancy, $|\Delta\varphi_{\mathrm{f}}|\sim(e^{S}/N)/\sqrt{N}$. Thus, as we expected, smaller bandwidths facilitate smaller deviations, which comes at a cost of more frequent, finite-sample-size-related misidentifications of the highest peak.